# Privacy-preserving Blockchain-enabled Parametric Insurance via Remote Sensing and IoT

### v2

@@ VIII-E Implementation of Bushfire Insurance by Sonic Protocol
  • Example 1: This problem is to decide if $w$ is binary: $w\in\{0,1\}$. This example can be represented by a single constraint equation: $w(w-1)=0$, equivalent to the following constraint system ${\mathscr{C}}$ with one multiplicative constraint and three linear constraints:
@@
    $({\mathbf{a}}_{1},{\mathbf{b}}_{1},{\mathbf{c}}_{1})$=$(w,w,w^{2})$ is satisfiable in ${\mathscr{C}}$, if $w\in\{0,1\}$.
@@
    $\mathbf{a}_{i}\cdot\mathbf{b}_{i}&=\mathbf{c}_{i},&\mbox{for\ }1\leq i\leq k\\ {\mathbf{a}}_{i}-{\mathbf{b}}_{i}&=0,&\mbox{for\ }1\leq i\leq k\\ {\mathbf{a}}_{i}2^{i-1}-{\mathbf{c}}_{i}&=0,&\mbox{for\ }1\leq i\leq k\\ \sum_{i=1}^{k}{\mathbf{c}}_{i}&=w\end{array}\right.\qquad(\mathscr{C})$
    $({\mathbf{a}}_{i},{\mathbf{b}}_{i},{\mathbf{c}}_{i})$=$(b_{i},b_{i},b_{i}^{2})$ is satisfiable in ${\mathscr{C}}$, if $w$ is non-negative and $(b_{i})_{i=1}^{k}$ represent the (scaled) bit-decomposition of $w$.
TABLE XV: Settings of vectors $({\bf a},{\bf b},{\bf c})$ for parametric bushfire insurance.
@@
$1,...,n$ $(\mathbf{r}^{-}_{i}-\mathbf{s}^{-}_{i})=\mathbf{n}^{-}_{i}(\mathbf{r}^{-}_{i}+\mathbf{s}^{-}_{i})+\mathbf{\theta}^{-}_{i},i\in\{1,...,n\}$ $(\mathbf{n}_{i}^{-})_{i=1}^{n}$ $(\mathbf{r}_{i}^{-}+\mathbf{s}_{i}^{-})_{i=1}^{n}$ $(\mathbf{r}_{i}^{-}-\mathbf{s}_{i}^{-}-\mathbf{\theta}_{i}^{-})_{i=1}^{n}$
$n+1,...,2n$ $(\mathbf{r}^{+}_{i}-\mathbf{s}^{+}_{i})=\mathbf{n}^{+}_{i}(\mathbf{r}^{+}_{i}+\mathbf{s}^{+}_{i})+\mathbf{\theta}^{+}_{i},i\in\{1,...,n\}$ $(\mathbf{n}_{i}^{+})_{i=1}^{n}$ $(\mathbf{r}_{i}^{+}+\mathbf{s}_{i}^{+})_{i=1}^{n}$ $(\mathbf{r}_{i}^{+}-\mathbf{s}_{i}^{+}-\mathbf{\theta}_{i}^{+})_{i=1}^{n}$
$2n+1,...,3n$ $\mathbf{i}_{i}(\mathbf{n}^{-}_{i}-\mathbf{n}^{+}_{i}-\kappa)=\sum_{j=1}^{k}\mathbf{e}_{i,j},i\in\{1,...,n\}$ $(\mathbf{i}_{i})_{i=1}^{n}$ $(\mathbf{n}_{i}^{-}-\mathbf{n}_{i}^{+}-\kappa)_{i=1}^{n}$ $(\sum_{j=1}^{k}\mathbf{e}_{i,j})_{i=1}^{n}$
$3n+1,...,4n$ $(\mathbf{\theta}_{i}^{-})^{2},i\in\{1,...,n\}$ $(\mathbf{\theta}_{i}^{-})_{i=1}^{n}$ $(\mathbf{\theta}_{i}^{-})_{i=1}^{n}$ ${(\mathbf{\theta}_{i}^{-})^{2}}_{i=1}^{n}$
$4n+1,...,5n$ $(\mathbf{\theta}_{i}^{+})^{2},i\in\{1,...,n\}$ $(\mathbf{\theta}_{i}^{+})_{i=1}^{n}$ $(\mathbf{\theta}_{i}^{+})_{i=1}^{n}$ ${(\mathbf{\theta}_{i}^{+})^{2}}_{i=1}^{n}$
$5n+1,...,6n$ $\mathbf{i}_{i}\in\{0,1\},i\in\{1,...,n\}$ $(\mathbf{i}_{i})_{i=1}^{n}$ $(1-\mathbf{i}_{i})_{i=1}^{n}$ $\mathbf{0}$
$6n+1,...,6n+nk$ $\mathbf{e}_{i}\geq 0$ $(\mathbf{e}_{i,j})_{i=1,j=1}^{n,k}$ $(\mathbf{e}_{i,j}-2^{j-1})_{i=1,j=1}^{n,k}$ $\mathbf{0}$
$6n+nk+1,...,6n+nk+k$ $\theta_{d}\geq 0$ $(\mathbf{e}_{\theta_{d},j})_{j=1}^{k}$ $(\mathbf{e}_{\theta_{d},j}-2^{j-1})_{j=1}^{k}$ $\mathbf{0}$
$6n+(n+1)k+1,...,6n+(n+2)k$ $G\geq 0$ $(\mathbf{e}_{G,j})_{j=1}^{k}$ $(\mathbf{e}_{G,j}-2^{j-1})_{j=1}^{k}$ $\mathbf{0}$
$6n+(n+2)k+1,...,7n+(n+2)k$ Input from data source $(\mathbf{r}_{i}^{-})_{i=1}^{n}$ $\mathbf{0}$ $\mathbf{0}$
$7n+(n+2)k+1,...,8n+(n+2)k$ Input from data source $(\mathbf{s}_{i}^{-})_{i=1}^{n}$ $\mathbf{0}$ $\mathbf{0}$
$8n+(n+2)k+1,...,9n+(n+2)k$ Input from data source $(\mathbf{r}_{i}^{+})_{i=1}^{n}$ $\mathbf{0}$ $\mathbf{0}$
$9n+(n+2)k+1,...,10n+(n+2)k$ Input from data source $(\mathbf{s}_{i}^{+})_{i=1}^{n}$ $\mathbf{0}$ $\mathbf{0}$
VIII-F Security of the Enhanced zk-SNARK Protocol
@@
  • Non-interactiveness: There are several interactive protocols with public coins, meaning that the verifier must keep its internal state public. These interactive protocols with public coins can be converted to non-interactive protocols by Fiat-Shamir heuristic [47]. We presented the interactive zk-SNARK protocol with public coins in Table VI and Table XIII, together with the modifications (using the Fiat-Shamir heuristic) to achieve non-interactive protocols in grey next to the interactive steps.
@@
    Perfects Honest-Verifier Zero Knowledge Assume an arbitrary polynomial-time simulator Sim who can access all the public input of the protocol and the SRS strings from the $J$ data providers. It chooses random vectors $\mathbf{a},\mathbf{b}$ from ${\mathbb{F}}_{p}$ of length $n$ and sets $\mathbf{c}=\mathbf{a}\cdot\mathbf{b}$. It then chooses $J$ random vectors $\mathbf{d}_{1},...,d_{J}$, of length $m_{j}$ for $j\in{1,...,J}$. Then the simulator computes $\tilde{r}[X,Y],d_{1}[X],...,d_{J}[X]$ and $r[X,Y],t[X,Y]$. Then, the simulator performs the same as the prover described in Table VI with respect to the polynomials. As described in Section V-A, the prover only reveals $5+2J$ evaluations of $r[X,Y]$, (including $r(z,1),r(zy,1),\tilde{r},g^{r(X,1)},g^{\tilde{r}(X,1)}$ and $d_{j}[z],g^{d_{j}},j\in\{1,...,J\}$). Therefore, we set $6+2j$ random blinders with random coefficients and powers from $-2n-1$ to $-2n-6-2j$. Therefore, for a verifier obtained less than $6+2j$ openings, the prover’s polynomials are indistinguishable from the random polynomials from the simulator. All other polynomials are either public input or computed from $r[X,Y]$, hence preserving perfect honest-verifier zero knowledge.

### v3

@@ VIII-G Implementation of Bushfire Insurance by Sonic Protocol
  • Example 1: This problem is to decide if $w$ is binary: $w\in\{0,1\}$. This example can be represented by a single constraint equation: w(w1)=0w(w-1)=0, equivalent to the following constraint system $\mathscr{C}$} with one multiplicative constraint and three linear constraints:
@@
    $({\mathbf{a}}_{1},{\mathbf{b}}_{1},{\mathbf{c}}_{1})$$(w,w,w^{2})$ is satisfiable in $\mathscr{C}$}, if $w\in\{0,1\}$.
@@
[…]
    $({\mathbf{a}}_{i},{\mathbf{b}}_{i},{\mathbf{c}}_{i})$$(b_{i},b_{i},b_{i}^{2})$ is satisfiable in $\mathscr{C}$}, if $w$ is non-negative and $(b_{i})_{i=1}^{k}$ represent the (scaled) bit-decomposition of $w$.
TABLE XVII: Settings of vectors $({\bf a},{\bf b},{\bf c})$ for parametric bushfire insurance.
@@
1,,n1,...,n $(\mathbf{r}^{-}_{i}-\mathbf{s}^{-}_{i})=\mathbf{n}^{-}_{i}(\mathbf{r}^{-}_{i}+\mathbf{s}^{-}_{i})+\mathbf{\theta}^{-}_{i},i\in\{1,...,n\}$ $(\mathbf{n}_{i}^{-})_{i=1}^{n}$ $(\mathbf{r}_{i}^{-}+\mathbf{s}_{i}^{-})_{i=1}^{n}$ $(\mathbf{r}_{i}^{-}-\mathbf{s}_{i}^{-}-\mathbf{\theta}_{i}^{-})_{i=1}^{n}$
n+1,,2nn+1,...,2n $(\mathbf{r}^{+}_{i}-\mathbf{s}^{+}_{i})=\mathbf{n}^{+}_{i}(\mathbf{r}^{+}_{i}+\mathbf{s}^{+}_{i})+\mathbf{\theta}^{+}_{i},i\in\{1,...,n\}$ $(\mathbf{n}_{i}^{+})_{i=1}^{n}$ $(\mathbf{r}_{i}^{+}+\mathbf{s}_{i}^{+})_{i=1}^{n}$ $(\mathbf{r}_{i}^{+}-\mathbf{s}_{i}^{+}-\mathbf{\theta}_{i}^{+})_{i=1}^{n}$
2n+1,,3n2n+1,...,3n $\mathbf{i}_{i}(\mathbf{n}^{-}_{i}-\mathbf{n}^{+}_{i}-\kappa)=\sum_{j=1}^{k}\mathbf{e}_{i,j},i\in\{1,...,n\}$ $(\mathbf{i}_{i})_{i=1}^{n}$ $(\mathbf{n}_{i}^{-}-\mathbf{n}_{i}^{+}-\kappa)_{i=1}^{n}$ $(\sum_{j=1}^{k}\mathbf{e}_{i,j})_{i=1}^{n}$
3n+1,,4n3n+1,...,4n $(\mathbf{\theta}_{i}^{-})^{2},i\in\{1,...,n\}$ $(\mathbf{\theta}_{i}^{-})_{i=1}^{n}$ $(\mathbf{\theta}_{i}^{-})_{i=1}^{n}$ ${(\mathbf{\theta}_{i}^{-})^{2}}_{i=1}^{n}$
4n+1,,5n4n+1,...,5n $(\mathbf{\theta}_{i}^{+})^{2},i\in\{1,...,n\}$ $(\mathbf{\theta}_{i}^{+})_{i=1}^{n}$ $(\mathbf{\theta}_{i}^{+})_{i=1}^{n}$ ${(\mathbf{\theta}_{i}^{+})^{2}}_{i=1}^{n}$
5n+1,,6n5n+1,...,6n $\mathbf{i}_{i}\in\{0,1\},i\in\{1,...,n\}$ $(\mathbf{i}_{i})_{i=1}^{n}$ $(1-\mathbf{i}_{i})_{i=1}^{n}$ $\mathbf{0}$
6n+1,,6n+nk6n+1,...,6n+nk $\mathbf{e}_{i}\geq 0$ $(\mathbf{e}_{i,j})_{i=1,j=1}^{n,k}$ $(\mathbf{e}_{i,j}-2^{j-1})_{i=1,j=1}^{n,k}$ $\mathbf{0}$
6n+nk+1,,6n+nk+k6n+nk+1,...,6n+nk+k $\theta_{d}\geq 0$ $(\mathbf{e}_{\theta_{d},j})_{j=1}^{k}$ $(\mathbf{e}_{\theta_{d},j}-2^{j-1})_{j=1}^{k}$ $\mathbf{0}$
6n+(n+1)k+1,,6n+(n+2)k6n+(n+1)k+1,...,6n+(n+2)k $G\geq 0$ $(\mathbf{e}_{G,j})_{j=1}^{k}$ $(\mathbf{e}_{G,j}-2^{j-1})_{j=1}^{k}$ $\mathbf{0}$
6n+(n+2)k+1,,7n+(n+2)k6n+(n+2)k+1,...,7n+(n+2)k Input from data source $(\mathbf{r}_{i}^{-})_{i=1}^{n}$ $\mathbf{0}$ $\mathbf{0}$
7n+(n+2)k+1,,8n+(n+2)k7n+(n+2)k+1,...,8n+(n+2)k Input from data source $(\mathbf{s}_{i}^{-})_{i=1}^{n}$ $\mathbf{0}$ $\mathbf{0}$
8n+(n+2)k+1,,9n+(n+2)k8n+(n+2)k+1,...,9n+(n+2)k Input from data source $(\mathbf{r}_{i}^{+})_{i=1}^{n}$ $\mathbf{0}$ $\mathbf{0}$
9n+(n+2)k+1,,10n+(n+2)k9n+(n+2)k+1,...,10n+(n+2)k Input from data source $(\mathbf{s}_{i}^{+})_{i=1}^{n}$ $\mathbf{0}$ $\mathbf{0}$
VIII-H Security of the Enhanced zk-SNARK Protocol
@@
  • Non-interactiveness: There are several interactive protocols with public coins, meaning that the verifier must keep its internal state public. These interactive protocols with public coins can be converted to non-interactive protocols by Fiat-Shamir heuristic [45]. We presented the interactive zk-SNARK protocol with public coins in Table VI and Table XV, together with the modifications (using the Fiat-Shamir heuristic) to achieve non-interactive protocols in grey next to the interactive steps.
@@
    Perfects Honest-Verifier Zero Knowledge Assume an arbitrary polynomial-time simulator Sim who can access all the public input of the protocol and the SRS strings from the $J$ data providers. It chooses random vectors $\mathbf{a},\mathbf{b}$ from ${\mathbb{F}}_{p}$ of length $n$ and sets $\mathbf{c}=\mathbf{a}\cdot\mathbf{b}$. It then chooses $J$ random vectors $\mathbf{d}_{1},...,d_{J}$, of length $m_{j}$ for $j\in{1,...,J}$. Then the simulator computes $\tilde{r}[X,Y],d_{1}[X],...,d_{J}[X]$ and $r[X,Y],t[X,Y]$. Then, the simulator performs the same as the prover described in Table VI with respect to the polynomials. As described in Section V-A, the prover only reveals $5+2J$ evaluations of $r[X,Y]$, (including $r(z,1),r(zy,1),\tilde{r},g^{r(X,1)},g^{\tilde{r}(X,1)}$ and $d_{j}[z],g^{d_{j}},j\in\{1,...,J\}$). Therefore, we set $6+2j$ random blinders with random coefficients and powers from 2n1-2n-1 to 2n62j-2n-6-2j. Therefore, for a verifier obtained less than $6+2j$ openings, the prover’s polynomials are indistinguishable from the random polynomials from the simulator. All other polynomials are either public input or computed from $r[X,Y]$, hence preserving perfect honest-verifier zero knowledge.